lasgow Enterprises started the period with 80 units in beginning inventory that cost $1.90 each. During the period, the company purchased inventory items as follows. Glasgow sold 265 units after purchase 3 for $1.90 each.

Purchase No. of Items Cost
1 400 $ 2.40
2 100 $ 2.50
3 60 $ 2.90

Glasgow's ending inventory under weighted average would be approximately: a. 900 b. 636 c. 786 d. 907
Answer #1

Total cost = (80*1.90)+(400*2.4)+(100*2.5)+(60*2.9)

= 1536

Total units = 80 + 400 +100 + 60 = 640

Average cost per unit = 1536/640 = 2.40

Ending inventory = Units available - Units sold = 640 - 265 = 375

Cost of ending inventory = 375*2.4 = 900

Option A